How much heat in kJ gets rid of condensing 7.13 g of water vapor at 100 ° C and then cools the water formed to 25 ° C?

 [ΔHvap = 40.7 kJ / mol at 100 ° C and the natural heat of water c (H2O (l)) = 4.18 J g-1 ° C-1]

Note. look up the units in the calculations and give the answer in kJ
Answer #1
We have 7.13 g of water vapor at 100 ° C 
 So, moles of water present = 7.13 g / (18 g/mol) = 0.396 moles
 Condensation of water vapour to water at 25°C has two steps :
(i) water vapour to water at 100 C
  heat lost is latent heat of vapourization , we have ΔHvap = 40.7kJ/mol  
 So, amount of heat lost during this step , 
   Q1 = 0.396 moles*40.7kJ/mol  = 16.12 kJ

(ii) It cools to 25 ° C
Heat is lost during ,
Q2 = m*ΔT* c   ( m = mass of water ;ΔT = temp. change (in degree) : c = specific/natural heat of water  = 4.18 J g-1°C-1 

Q2 = 7.13 g* 75 oC *  4.18 J g-1°C-1  = 2235.2 J = 2.24 kJ

So, total amount of heat gets rid = Q1 + Q2 = 18.36 kJ